OpenStudy (anonymous):

If pq=3x, sq=5x-8, and os=x+1, find a.pq b.sq c.os d.oq * i will draw the shape

OpenStudy (anonymous):



OpenStudy (anonymous):

PA and SQ are tangents to the circle?

OpenStudy (phi):

PQ = SQ so you can solve for x: 3x= 5x-8

OpenStudy (phi):

@dpaInc presumably. I giving him the benefit of the doubt.

OpenStudy (anonymous):

***PQ

OpenStudy (phi):

The only "tricky" part is knowing that the tangent and radius form a right angle, so you can use pythagoras to find QO, then subtract the radius OP to get QR

OpenStudy (campbell_st):

pq = sq tangents from an external point are equal then pq = sq 3x = 5x - 8 8 = 2x x = 4 pq = 12, sq = 12

OpenStudy (campbell_st):

os = 4 + 1 = 5

OpenStudy (campbell_st):

and to find oq use pythagoras' theorem the tangent and radius intersect at right angles... so \[oq^2 = 12^2 + 5^2\] oq = 13

OpenStudy (anonymous):

thank you guys.